# Instruct-ReID: A Multi-purpose Person Re-identification Task with Instructions

### v4

@@ 4.3 Adaptive Triplet Loss
where $N_{tri\dagger}$ and $m$ denote the number of triplets and a hyper-parameter for the maximal margin, respectively. $d$ is a Euclidean distance function, i.e., $d(\mathbf{F}^{a}_{i},\mathbf{F}^{r}_{i})=\left|\left|\mathbf{F}^{a}_{i}-\mathbf{F}^{r}_{i}\right|\right|_{2}^{2}$. $\beta_{1}$ and $\beta_{2}$ are relatednesses between the anchor image and the corresponding reference image that take into account the identity consistency and the instruction similarity for the adaptive margin. Mathematically, they can be defined as
@@ 4.4 Overall Loss Function
where, $\mathcal{L}_{match}({F_{out}})=\mathcal{C}_{e}(\hat{y},F_{out})=\mathcal{C}_{e}[\hat{y},\mathcal{E}_{f}(\mathbf{F},\mathbf{F_{T}})]$, respectively. $\mathcal{C}_{e}$ represents a binary cross-entropy function, $\hat{y}$ is a 2-dimension one-hot vector representing the ground-truth label i.e., $[0,1]^{\intercal}$ for the positive pair and $[1,0]^{\intercal}$ for the negative pair, which is formed by matching text features with corresponding image features before inputting into the attention module $\mathcal{E}_{f}$.
@@ 5.1 Experimental Setups
Table 2: The performance of Clothse-Changing ReID of our method and the state-of-the-art methods. Mean average precision (mAP) and Top1 are used to quantify the accuracy. † denotes that the model is trained with multiple datasets. * denotes that the model is pre-trained on 4 million images.
@@ 5.2 Experimental Results
Clothes-Changing ReID (CC-ReID). As shown in Tab 2, IRM outperforms all state-of-the-art methods on LTCC, PRCC and VC-Clothes, showing that the model can effectively extract clothes-invariant features following the instructions, e.g., “Ignore clothes”. Specifically, on STL, IRM improves CAL [12], TransReID [14] by +5.9% mAP and +8.3% mAP on LTCC and VC-Clothes, respectively. Using multi-task training further improves the performance of IRM to 52.0% mAP on LTCC, and to reach a new state-of-the-art on PRCC with a 52.3% mAP. While multi-task learning leads to slightly lower performance than single-task learning on VC-Clothes, our method still achieves a higher Top-1 than TransReID. We conjecture that this drop is due to the domain gap between VC-Clothes (Synthetic) and other datasets (Real) and leave it for future work.
@@ 5.3 Ablation Study
Adaptive Triplet Loss. Tab 5 shows that adaptive triplet loss in MTL outperforms the traditional triplet loss by +0.3% mAP on average, indicating that the proposed loss boosts the model to learn more discriminative features following different instructions. In the STL scenario, for CC-ReID and Trad-ReID, the instructions are fixed sentences leading to the same performance of adaptive/traditional triplet loss. However, in the case of CTCC-ReID and LI-ReID, where instructions vary among samples, using adaptive triplet loss brings about +0.7%, +0.5% mAP gain, which shows the effectiveness of adaptive triplet loss in learning both identity and instruction similarity.
@@ C.1 Pedestrian Attribute Generation
coat color
”black coat”, ”blue coat”, ”gray coat”, ”green coat”, ”purple coat”, ”red coat”, ”white coat”, ”yellow coat”
trousers color
”black trousers”, ”blue trousers”, ”gray trousers”, ”green trousers”, ”purple trousers”, ”red trousers”, ”white trousers”, ”yellow trousers”
coat length
”agnostic length coat”, ”long sleeve coat”, ”short sleeve coat”, ” bareback coat”
trousers length
”shorts trousers”, ”skirt”, ”trousers”
gender code
”female”, ”agnostic gender”, ”male”
glass style
”without glasses”, ”with glasses”, ”with sunglasses”
hair color
”black hair”, ”agnostic color hair”, ”white hair”, ”yellow hair”
hair style
”bald hair”, ”agnostic style hair”, ”long hair”, ”short hair”
bag style
”backpack”, ”hand bag”, ”shoulder bag”, ”waist pack”, ”trolley”, ”agnostic style bag”, ”without bag”
cap style
”with hat”, ”without hat”
shoes color
”black shoes”, ”blue shoes”, ”gray shoes”, ”green shoes”, ”purple shoes”, ”red shoes”, ”white shoes”, ”yellow shoes”
shoes style
”boots”, ”leather shoes”, ”sandal”, ”walking shoes”
age
”adult”, ”child”, ”old”
person angle
”back”, ”front”, ”side”
pose
”lie”, ”pose agnostic”, ”sit”, ”stand”, ”stoop”
coat style
”business suit”, ”agnostic style coat”, ”dress”, ”jacket”, ”long coat”, ”shirt”, ”sweater”, ”t-shirt”
glove
”with glove”, ”agnostic glove”, ”without glove”
smoking
”smoking”, ”agnostic smoking”, ”without smoking”
umbrella
”with umbrella”, ”without umbrella”
uniform
”chef uniform”, ”common clothing”, ”firefighter uniform”, ”medical uniform”, ”office uniform”, ”agnostic uniform”, ”worker uniform”
@@ D.1 Instructions on traditional ReID, cloth-changing ReID and visual-infrared ReID
Trad-ReID
”do not change clothes”, ”maintain consistent clothes”, ”keep original clothes”, ”preserve current clothes”, ”retain existing clothes”, ”wear the same clothes”, ”stick with your clothes”, ”don’t alter your clothes”, ”no changes to clothes”, ”unchanged outfit”, ”clothes remain constant”, ”no clothing adjustments”, ”steady clothing choice”, ”clothing remains unchanged”, ”consistent clothing selection”, ”retain your clothing style”, ”clothing choice remains”, ”don’t swap clothes”, ”maintain clothing selection”, ”clothes stay the same”
CC-ReID
”change your clothes”, ”swap outfits”, ”switch attire”, ”get into a different outfit”, ”try on something new”, ”put on fresh clothing”, ”dress in alternative attire”, ”alter your outfit”, ”wear something else”, ”don a different ensemble”, ”trade your garments”, ”shift your wardrobe”, ”exchange your clothing”, ”update your attire”, ”replace your outfit”, ”clothe yourself differently”, ”switch your style”, ”update your look”, ”put on a new wardrobe”, ”ignore clothes”
VI-ReID
”retrieve cross-modality images”, ”fetch images across different modalities”, ”collect images from various modalities”, ”obtain images spanning different modalities”, ”retrieve images from diverse modalities”, ”gather images across modalities”, ”access images across different modalities”, ”acquire images spanning various modalities”, ”extract images from different modalities”, ”retrieve images across multiple modalities”, ”fetch images from distinct modalities”, ”collect images across various modalities”, ”access images from different modalities”, ”obtain images from diverse modalities”, ”gather images spanning different modalities”, ”extract images from various modalities”, ”retrieve images across varied modalities”, ”obtain images from distinct modalities”, ”access images across multiple modalities”, ”collect images spanning diverse modalities”
@@ E.1 Dataset Statistics of OmniReID
OmniReID Agreement This Agreement outlines the terms and conditions governing the use of OmniReID. By signing this agreement, the Recipient agrees to the following terms: The Recipient agrees the OmniReID does not claim Copyrights of Market1501, CUHK03, MSMT17, PRCC, VC-Clothes, LTCC, COCAS+, LLCM, CUHK-PEDES, SYNTH-PEDES and they SHOULD obtain these public datasets from these data providers. The Recipient agrees they must comply with ALL LICENSEs of Market1501, CUHK03, MSMT17, PRCC, VC-Clothes, LTCC, COCAS+, LLCM, CUHK-PEDES, SYNTH-PEDES when they use OmniReID. The Recipient agrees the demographic makeup of OmniReID is not representative of the broader population. The Recipient agrees OmniReID should only be available for non-commercial research purposes. Any other use, in particular any use for commercial purposes, is prohibited. The Recipient agrees not to use the data for any unlawful, unethical, or malicious purposes. The Recipient agrees not to further copy, publish or distribute any portion of the OmniReID. The Recipient agrees [OUR INSTITUTE] reserves the right to terminate the access to the OmniReID at any time. Name:
Organization/Affiliation:
Position:
Email:
Address:
Address (Line2):
City: Country:
Signature:
Date:

### v5

@@ 4.3 Adaptive Triplet Loss
where $N_{tri\dagger}$ and $m$ denote the number of triplets and a hyper-parameter for the maximal margin, respectively. $d$ is a Euclidean distance function, i.e., $d(\mathbf{F}^{a}_{i},\mathbf{F}^{r}_{i})=\left|\left|\mathbf{F}^{a}_{i}-\mathbf{F}^{r}_{i}\right|\right|_{2}^{2}$. $\beta_{1}$ and $\beta_{2}$ are relatednesses between the anchor image and the corresponding reference image that consider the identity consistency and instruction similarity for the adaptive margin. Mathematically, they are defined as
@@ 4.4 Overall Loss Function
where, $\mathcal{L}_{match}({F_{out}})=\mathcal{C}_{e}(\hat{y},F_{out})=\mathcal{C}_{e}[\hat{y},\mathcal{E}_{f}(\mathbf{F},\mathbf{F_{T}})]$, respectively. $\mathcal{C}_{e}$ is a binary cross-entropy loss, $\hat{y}$ is a 2-dimension one-hot vector representing the ground-truth label i.e., $[0,1]^{\intercal}$ for positive pair and $[1,0]^{\intercal}$ for negative pair, formed by matching text features with corresponding image features before inputting into the attention module $\mathcal{E}_{f}$.
@@ 5.1 Experimental Setups
Table 2: The performance of Clothes-Changing ReID of our method and the state-of-the-art methods. Mean average precision (mAP) and Top1 are used to quantify the accuracy. † denotes that the model is trained with multiple datasets. * denotes that the model is pre-trained on 4 million images.
@@ 5.2 Experimental Results
Clothes-Changing ReID (CC-ReID). As shown in Tab 2, IRM outperforms all state-of-the-art methods on LTCC, PRCC and VC-Clothes, showing that the model can effectively extract clothes-invariant features following the instructions, e.g., “Ignore clothes”. Specifically, on STL, IRM improves CAL [13], TransReID [16] by +5.9% mAP and +8.3% mAP on LTCC and VC-Clothes, respectively. On MTL, IRM further improves the performance on LTCC to 52.0% mAP and reaches a new state-of-the-art on PRCC with 52.3% mAP. Other methods like ACID [62], CCFA [14], AIM [61] and DCR-ReID [9] provide results of 384$\times$192 image size, we present more results based on this resolution of IRM in supplementary. While multi-task learning leads to slightly lower performance than single-task learning on VC-Clothes, our method still achieves a higher Top-1 than TransReID. We conjecture that this drop is due to the domain gap between VC-Clothes (Synthetic) and other datasets (Real) and leave it for future work.
@@ 5.3 Ablation Study
Adaptive Triplet Loss. Tab 5 shows that adaptive triplet loss in MTL outperforms the traditional triplet loss by +0.3% mAP on average, indicating that the proposed loss boosts the model to learn more discriminative features following different instructions. On STL, for CC-ReID and Trad-ReID, the instructions are fixed sentences leading to the same performance of adaptive/traditional triplet loss. However, in the case of CTCC-ReID and LI-ReID, where instructions vary among samples, using adaptive triplet loss brings about +0.7%, +0.5% mAP gain, which shows the effectiveness of adaptive triplet loss in learning both identity and instruction similarity.
@@ 6 Conclusion
Acknowledgement.This paper was supported by the National Natural Science Foundation of China (No.62127803), Key R&D Project of Zhejiang Province (No.2022C01056).
@@ C.1 Pedestrian Attribute Generation
coat color ”black coat”, ”blue coat”, ”gray coat”, ”green coat”, ”purple coat”, ”red coat”, ”white coat”, ”yellow coat”
trousers color ”black trousers”, ”blue trousers”, ”gray trousers”, ”green trousers”, ”purple trousers”, ”red trousers”, ”white trousers”, ”yellow trousers”
coat length ”agnostic length coat”, ”long sleeve coat”, ”short sleeve coat”, ” bareback coat”
trousers length ”shorts trousers”, ”skirt”, ”trousers”
gender code ”female”, ”agnostic gender”, ”male”
glass style ”without glasses”, ”with glasses”, ”with sunglasses”
hair color ”black hair”, ”agnostic color hair”, ”white hair”, ”yellow hair”
hair style ”bald hair”, ”agnostic style hair”, ”long hair”, ”short hair”
bag style ”backpack”, ”hand bag”, ”shoulder bag”, ”waist pack”, ”trolley”, ”agnostic style bag”, ”without bag”
cap style ”with hat”, ”without hat”
shoes color ”black shoes”, ”blue shoes”, ”gray shoes”, ”green shoes”, ”purple shoes”, ”red shoes”, ”white shoes”, ”yellow shoes”
shoes style ”boots”, ”leather shoes”, ”sandal”, ”walking shoes”
age ”adult”, ”child”, ”old”
person angle ”back”, ”front”, ”side”
pose ”lie”, ”pose agnostic”, ”sit”, ”stand”, ”stoop”
coat style ”business suit”, ”agnostic style coat”, ”dress”, ”jacket”, ”long coat”, ”shirt”, ”sweater”, ”t-shirt”
glove ”with glove”, ”agnostic glove”, ”without glove”
smoking ”smoking”, ”agnostic smoking”, ”without smoking”
umbrella ”with umbrella”, ”without umbrella”
uniform ”chef uniform”, ”common clothing”, ”firefighter uniform”, ”medical uniform”, ”office uniform”, ”agnostic uniform”, ”worker uniform”
@@ D.1 Instructions on traditional ReID, cloth-changing ReID and visual-infrared ReID
Trad-ReID ”do not change clothes”, ”maintain consistent clothes”, ”keep original clothes”, ”preserve current clothes”, ”retain existing clothes”, ”wear the same clothes”, ”stick with your clothes”, ”don’t alter your clothes”, ”no changes to clothes”, ”unchanged outfit”, ”clothes remain constant”, ”no clothing adjustments”, ”steady clothing choice”, ”clothing remains unchanged”, ”consistent clothing selection”, ”retain your clothing style”, ”clothing choice remains”, ”don’t swap clothes”, ”maintain clothing selection”, ”clothes stay the same”
CC-ReID ”change your clothes”, ”swap outfits”, ”switch attire”, ”get into a different outfit”, ”try on something new”, ”put on fresh clothing”, ”dress in alternative attire”, ”alter your outfit”, ”wear something else”, ”don a different ensemble”, ”trade your garments”, ”shift your wardrobe”, ”exchange your clothing”, ”update your attire”, ”replace your outfit”, ”clothe yourself differently”, ”switch your style”, ”update your look”, ”put on a new wardrobe”, ”ignore clothes”
VI-ReID ”retrieve cross-modality images”, ”fetch images across different modalities”, ”collect images from various modalities”, ”obtain images spanning different modalities”, ”retrieve images from diverse modalities”, ”gather images across modalities”, ”access images across different modalities”, ”acquire images spanning various modalities”, ”extract images from different modalities”, ”retrieve images across multiple modalities”, ”fetch images from distinct modalities”, ”collect images across various modalities”, ”access images from different modalities”, ”obtain images from diverse modalities”, ”gather images spanning different modalities”, ”extract images from various modalities”, ”retrieve images across varied modalities”, ”obtain images from distinct modalities”, ”access images across multiple modalities”, ”collect images spanning diverse modalities”
@@ E.2 Discussion of Ethical Issues
{mdframed}
[backgroundcolor=gray!10]
OmniReID Agreement
This Agreement outlines the terms and conditions governing the use of OmniReID. By signing this agreement, the Recipient agrees to the following terms:
  • The Recipient agrees the OmniReID does not claim Copyrights of Market1501, CUHK03, MSMT17, PRCC, VC-Clothes, LTCC, COCAS+, LLCM, CUHK-PEDES, SYNTH-PEDES and they SHOULD obtain these public datasets from these data providers.
  • The Recipient agrees they must comply with ALL LICENSEs of Market1501, CUHK03, MSMT17, PRCC, VC-Clothes, LTCC, COCAS+, LLCM, CUHK-PEDES, SYNTH-PEDES when they use OmniReID.
  • The Recipient agrees the demographic makeup of OmniReID is not representative of the broader population.
  • The Recipient agrees OmniReID should only be available for non-commercial research purposes. Any other use, in particular any use for commercial purposes, is prohibited.
  • The Recipient agrees not to use the data for any unlawful, unethical, or malicious purposes.
  • The Recipient agrees not to further copy, publish or distribute any portion of the OmniReID.
  • The Recipient agrees [OUR INSTITUTE] reserves the right to terminate the access to the OmniReID at any time.
    Name:
    Organization/Affiliation:
    Position:
    Email:
    Address:
    Address (Line2):
    City: Country:
    Signature:
    Date: